OpenStudy (anonymous):

Help asap please! 2. Jupiter has a radius of about 11.0 earth radii, and a mass 318 times the earth’s mass. Estimate the gravitational field on the surface of Jupiter compared to that on the earth. Show your work.

OpenStudy (anonymous):

F= G (m1)(m2)/r^2, where G is the gravitational constant, m1 is any mass (say, yours), and m2 is the mass of the planet. r is the radius. On Earth, Fearth = G(100kg)(Mearth)/r^2 On Jupiter, Fjupiter = G(100kg)(Mjupiter)/r^2 G, and m1 will be the same on both worlds, so: Fjupiter = Mjupiter / r^2 = 318 / 11^2 = 318 / 121 = 2.63 times as much on Earth,